OpenStudy (anonymous):

the instantaneous rate of change of y=f(x) with respect to x at x=2

OpenStudy (anonymous):

What is f(x)?

OpenStudy (anonymous):

f(x)=3x^2

OpenStudy (anonymous):

Good deal. Take the derivative of f(x) and then evaluate it at x=2, since instantaneous rate of change is just code for derivative. For f(x)=3x^2, f'(x) = 6x, so at x=2, the derivative is 12.

OpenStudy (anonymous):

yeah i have the slope or rather the deritive, which infact is 12 smart,

OpenStudy (anonymous):

f(x)=3x^2 f'(x)=6(x) f'(2)=6*2=12

OpenStudy (anonymous):

I'm just not sure, the definition of the instananeous rate of change is the lim as h -> 0 of the difference quotient, my professor did not explain this well

OpenStudy (anonymous):

The lim as h -> 0 is the definition of the derivative. The derivative is the collecting instantaneous rates of change for the function f(x). So f'(x) at a point is that points instantaneous rate of change. Shanna and etothepiiplusone are correct.

OpenStudy (anonymous):

thanks :)